Given: f(x) = 5x – 1; Find f(a-2)
Answer by Edwin McCravy(20086)


Given: f(x) = 5x – 1; Find f(a-2)

       f(x) = 5x - 1

Just replace every x you see with (a-2)

       f((a-2)) = 5(a-2) - 1

then simplify

       f(a-2) = 5a - 10 - 1

       f(a-2) = 5a - 11

That's all there is to it!

Edwin